OpenStudy (anonymous):

Simplify a^(3 loga 5) given that a>1

hartnn (hartnn):

\(m \log n = \log n^m\) so, what about \(3 \log_a5 =...?\)

OpenStudy (anonymous):

ans is 5^3 =125

OpenStudy (anonymous):

@hartnn I understand it would be loga 125

OpenStudy (anonymous):

but I'm not sure what to do next :S

OpenStudy (isaiah.feynman):